OpenStudy (rainbow_rocks03):

PLZ help will MEDAL+FAN for BEST response. A bag contains 6 green marbles, 18 red marbles, 2 white marbles, and 4 black marbles. What is the probability of drawing a black marble out of the bag?

OpenStudy (rainbow_rocks03):

No direct answers give the explanation too.

OpenStudy (rainbow_rocks03):

A. 1/3 B. 1/30 C. 2/15 D. 1/4

OpenStudy (rainbow_rocks03):

@samjordon @NeonStrawsForever @TheSmartOne @sleepyjess

OpenStudy (kropot72):

\[\large P(black)=\frac{number\ of\ black}{total\ number\ of\ marbles}\]

OpenStudy (rainbow_rocks03):

4 black/30 total

OpenStudy (kropot72):

Correct. Now simplify 4/30.

OpenStudy (rainbow_rocks03):

How do I simplify it? I divide 4 on both sides?

OpenStudy (kropot72):

The highest common factor of 4 and 30 is 2. So divide 4 by 2 and divide 30 by 2.

OpenStudy (anonymous):

c

OpenStudy (rainbow_rocks03):

oh ok. 4 / 2 = 2 & 30 / 2 = 15 so it is 2/15

OpenStudy (kropot72):

Correct.

OpenStudy (rainbow_rocks03):

No direct answers I aready typed that. :)

OpenStudy (rainbow_rocks03):

thx for the help @kropot72

OpenStudy (kropot72):

You're welcome :)